OpenStudy (anonymous):

The equation y'' + y' -2y = x^2 is called a differential equation because it involves an unknown function y and its derivatives y' and y''. Find constants A,B, and C such that the function y=Ax^2+Bx+c satisfies the equation

OpenStudy (anonymous):

If y=Ax^2 + Bx + C y' = 2Ax + B y''= 1 Is what I have so far

OpenStudy (anonymous):

BTW, this is a calculus I class

OpenStudy (turingtest):

substitute those expressions for y, y', and y'' into the equation above

OpenStudy (turingtest):

tell me what you get, or if you need more explanation

OpenStudy (anonymous):

Y=AX^2+BX+C Y'=2AX+B Y''=2A Y''+Y'-2Y=2A+2AX+B-2AX^2-2BX-2C=X^2, then -2AX^2+(2A-2B)X+B-2C=X^2,and -2A=1, and A=-1/2 2A-2B=0, then B=A=-1/2 B-2C=0, then C=B/2=-1/4